A is 27 years old. B is one third older. How many years ago was the ratio of A's age to B's age equal to 4:7?

Answer by checkley79(3341)

4/7=(A-X)/(B-X)
4/7=(27-X)/(27+9-X) CROSS MULTIPLY.
7(27-X)=4(36-X)
189-7X=144-4X
-7X+4X=144-189
-3X=-45
X=-45/-3
X=15 YEARS AGE THEIR AGES WERE 4:7
PROOF:
(27-15)/(36-15)
12/21=4/7
12*7=21*4
84=84